OpenStudy (anonymous):

Find the number of sides of a regular polygon if the sum of the measures of the interior angles is 900°. The number of sides is:

OpenStudy (anonymous):

sum of interior angles = 180*(n-2) where n denotes the number of sides

OpenStudy (anonymous):

n-2 = 900/180 or n= 2+ (900/180)

OpenStudy (anonymous):

so now you can find n

OpenStudy (anonymous):

thank U:)

OpenStudy (anonymous):

yw